OpenStudy (pphalke):

In a triangle, the measure of the first angle is three times the measure of the second angle. The measure of the third angle is 55 degrees more than the measure of the second angle. use the fact that the sum of the measures of three angles of a triangle is 180 degrees to find the measures of each angle.

Vocaloid (vocaloid):

let the second angle = x first angle = 3x third angle = x + 55 therefore x + 3x + x + 55 = 180 solve for x

Vocaloid (vocaloid):

still there?

OpenStudy (pphalke):

yes sorry slow computer

Vocaloid (vocaloid):

no problem, do you know how to solve for x?

OpenStudy (pphalke):

no I don't

Vocaloid (vocaloid):

x + 3x + x = ?

OpenStudy (pphalke):

5x

Vocaloid (vocaloid):

good now, 5x + 55 = 180 subtract 55 from both sides. what do you get?

OpenStudy (pphalke):

125

Vocaloid (vocaloid):

ok, now 5x = 125 what does x = ?

OpenStudy (pphalke):

25

Vocaloid (vocaloid):

good! so the second angle is 25 now we need to figure out the other two angles we said before that the first angle is 3 times the second angle so what is the first angle?

OpenStudy (pphalke):

I don't know

Vocaloid (vocaloid):

think carefully. second angle = 25 first angle = 3 * second angle = _____?

Vocaloid (vocaloid):

* means multiplication

OpenStudy (pphalke):

25 *3=75

Vocaloid (vocaloid):

good! second angle = 25 third angle = second angle + 55 = ???

OpenStudy (pphalke):

80

Vocaloid (vocaloid):

very good! so your angles are 75, 80, and 25 in that order. and that's it!

OpenStudy (pphalke):

first angle 75 second angle 25 and third angle 80?

Vocaloid (vocaloid):

oh right, sorry, you're right ^^

Vocaloid (vocaloid):

second angle is 25